# Verisimilitude as Boon and Bane: How People Initiate Opportunistic Interactions at Professional Events in Social VR

### v1

@@ Abstract.
social VR, opportunistic interaction, professional events, non-verbal cues
$c$s: Human-centered computing Collaborative and social computingccs: Human-centered computing Collaborative and social computing theory, concepts and paradigms
@@ 1. Introduction
The current research connects social VR with opportunistic interaction in professional settings—the unstructured yet essential exchanges that enable social bonding and knowledge transfer among members of the same workspace or community. Previous work in CSCW and communication research has delved deeply into each of these areas. Behavioral studies on how people coordinate spontaneity in workplace communication can date back to early 1990s, followed by two decades of experimental work aiming to reproduce this coordination in remote settings (e.g.,  (Tang and Rua, 1994; Fish et al., 1990; Isaacs and Tang, 1993)). Under a separate endeavor, recent scholarship has explored the emerging use of social VR for social hangouts and networking in informal meetups (e.g.,  (McVeigh-Schultz et al., 2019; Moustafa and Steed, 2018; Maloney and Freeman, 2020)). Little research, however, exists at the intersection of these two, although social VR arguably holds greater promise than most other mediums for supporting high-stakes interactions.
We offer an in-depth qualitative account to address this oversight. Specifically, we investigate factors that contribute to the successes and failures of opportunistic interaction in VR-mediated professional events. All participants involved in this investigation were authentic users of social VR. With their consent, we shadowed each person during one or more public professional events they have signed up for, closely observing the way they navigated each opportunistic interaction in its social context. We also conducted post-event interviews with the person to gather their retrospections, enabling triangulation. The data obtained from these steps allow us to develop a nuanced understanding of how interaction unfolds in a context jointly characterized by spontaneity, professionalism, and virtuality.
Our findings provide rich empirical insights that solidify the promise and current roadblocks toward effectively initiating opportunistic interaction in social VR. At the center of these findings lies the notion of verisimilitude, or the participant’s perceived extent to which a social VR environment has preserved the real-world meanings of various cues exchanged there. Our participants’ behaviors and interview responses revealed three distinct steps leading to the eventual occurrence of opportunistic interaction in professional settings: availability recognition, attention capture, and ice-breaking. At each step, an individual’s assessment of verisimilitude could be disrupted by a varied set of issues. Participants demonstrated an escalating need for high verisimilitude as they advanced through the three steps. They showed an increasing reliance on the host’s assistance when this need was not met. These insights provide actionable guidance for the design of future social VR platforms, as well as for people who navigate professional interactions via those platforms.
@@ 2. Related Work
In this section, we begin by reviewing literature on opportunistic interactions in traditional face-to-face workplace contexts, emphasizing its crucial role in professional settings and how it often begins as a momentary interaction using both verbal greetings and non-verbal cues (Section 2.1). We then examine how opportunistic communication unfolds and is facilitated through virtual video-conferencing, comparing the similarities and differences between face-to-face and online video-mediated versions of opportunistic interactions for professional purposes (Section 2.2). Next, we consider social VR as a platform for opportunistic interactions, noting that much of the existing research explores VR social interactions in non-professional contexts. We discuss the challenges and strategies people use in these settings (Section 2.3) and conclude with insights on why these strategies may not directly apply to professional contexts.
2.1. Opportunistic Interactions in Face-to-Face Professional Settings
@@
Opportunistic interactions play a crucial role in professional, face-to-face settings (Kraut et al., 1990; Fay, 2011). Studies show that in office environments, opportunistic interactions can constitute up to approximately 92% of all conversations (Whittaker et al., 1994), fostering strong social bonds among colleagues (Holmes and Marra, 2004), enabling spontaneous problem-solving without the formal pressures of meetings (Conrad and Poole, 1997), and helping to build meaningful interpersonal relationships (Duck, 1993). Beyond the workplace, opportunistic interactions is equally essential in professional contexts like conferences and workshops, where it facilitates networking, secures business clients and contracts, and promotes academic collaborations (Oester et al., 2017; McCarthy et al., 2004).
@@
Despite the prevalence of opportunistic interactions in professional settings, initiating these interactions can feel challenging due to the complex and dynamic behaviors involved (Pillet-Shore, 2018). Research in communication and psychology has therefore examined how face-to-face conversations are initiated, identifying two main features: audible and non-verbal cues (Kendon and Ferber, 1973; Pillet-Shore, 2012). For example, Pillet-Shore analyzed 337 video-recorded openings of face-to-face interactions, finding that people often use greeting utterances like Hello or Good morning to acknowledge another’s presence as a conversation starter. While these greetings may appear similar, they carry subtle nuances along a prosodic continuum, with ‘large’ greetings expressing positivity and ‘small’ ones conveying a neutral stance (Pillet-Shore, 2012).
@@
Non-verbal signals are also essential and often work in combination with audible cues to initiate opportunistic interactions in face-to-face settings. Previous research observed that visible smiling frequently accompanies greeting utterances, facilitating conversation openings (Kendon and Ferber, 1973). Additionally, individuals may adjust their body positions during moments of co-presence to ease the initiation process, allowing them to reset spatial and territorial orientations for a smoother approach (Goodwin, 1981; Mortensen and Hazel, 2014). In another study, Keevallik analyzed a full-day recording of 11 young people working together, finding that factors such as posture, gaze, and the bodily orientation of fellow workers contribute to attracting responses from others (Keevallik, 2018).
@@
In short, the above literature reveals that multiple elements, both verbal and non-verbal, contribute to the successful initiation of opportunistic interactions face-to-face. It is worth noting that, while multiple signals influence a dyad’s perceived availability and willingness to engage, the duration of this initial stage is typically as brief as a few seconds (Pillet-Shore, 2018). Although turn-taking generally becomes the norm once the conversation has started (Stivers et al., 2009; Wiemann and Knapp, 2017), in face-to-face interaction, the initiation phase can be viewed as a single, cohesive, and brief process.
2.2. Computer-Mediated Informal Communication in Professional Contexts
@@
Over recent decades, advancements in computer-mediated communication (CMC) tools, including instant messaging and video conferencing, have profoundly reshaped communication in professional contexts, transitioning from solely co-located, face-to-face interactions to more flexible, frequently remote environments (Fish et al., 1992; Zhao and Rosson, 2009; Egido, 1988). The shift was accelerated by the COVID-19 pandemic, which prompted many companies to adopt fully remote work models through team-based video conferencing (Fana et al., 2020; Bennett et al., 2021; Hacker et al., 2020). Despite this flexibility, remote settings greatly limit opportunistic interactions that naturally occur in physical workspaces (Fay, 2011; Bleakley et al., 2022). Previous literature attributes this reduction in CMC-based opportunistic interactions to two main factors: the lack of spatial relationships in video conferencing—similar to hallways in physical offices that facilitate casual encounters (Rudnicka et al., 2022; Isaacs and Tang, 1993)—and the absence of rich cues prevalent in face-to-face interactions (Grayson and Monk, 2003; Fish et al., 1992; Doherty-Sneddon et al., 1997).
@@
Regarding the first challenge, CSCW researchers have long investigated ways to foster a sense of co-location in distributed video conferencing (e.g.,  (Hauber et al., 2006; Raskar et al., 1998; Gibbs et al., 1999; Vertegaal, 1999; Benford et al., 2001; Elrod et al., 1992)). For example, early work by Fish et al. attempted to simulate physical proximity by placing human-scale screens in workspaces on different floors, encouraging workers in separate locations to engage in spontaneous conversations (Fish et al., 1990). Later, CRUISER (Fish et al., 1992; Root, 1988) and Montage (Tang et al., 1994; Tang and Rua, 1994) extended this concept to desktop-based video software. CRUISER was inspired by the concept of roaming a hallway and meeting people opportunistically; it facilitated both intentional and system-generated spontaneous video connections between users. Montage, on the other hand, introduced a gradual appearance and fading effect for the other person’s image, simulating movement closer to or farther away to mimic interactions in physical spaces. While these early explorations were not directly integrated into later video conferencing systems, the concept of rendering proximity has become widely adopted. For example, commercial platforms like Gather.town  [^1] Gather.town: https://www.gather.town/ use a shared 2D spatial layout where distributed users can see and approach each other. When users come into close proximity, a traditional video chat interface appears, allowing them to initiate spontaneous interactions. A recent study by Sanchez et al. found that such features indeed enhanced users’ awareness of each other, making it easier to determine when to interact, interrupt, or initiate a conversation. However, Gather.town users still reported insufficient information on whether others were actually interruptible for a conversation, essentially due to the lack of non-verbal cues within the 2D spatial layout interfaces (Palos-Sanchez et al., 2023).
@@
It should be noted that non-verbal cues such as facial expressions, gaze, and gestures are widely utilized during video conferencing (Nguyen and Canny, 2009). For example, studies have shown that gestures frequently enhance verbal communication, posture and facial expressions convey subtle emotions (Isaacs and Tang, 1993), whereas gaze can foster trust when well-perceived through video streaming (Nguyen and Canny, 2007). However, unlike in face-to-face opportunistic interactions, these non-verbal cues generally only come into play after a conversation has started, as the basic design of video conferencing does not enable their use prior to video streaming. Additionally, research has suggested that in video conferencing, non-verbal cues can create an illusion of transparency, leading users to overestimate their effectiveness (J. Sears et al., 2013).
In summary, CMC tools have become essential in today’s professional environments, enabling remote and flexible work arrangements. However, the reliance on video cameras and 2D screens weakens the sense of co-presence and limits the effectiveness of non-verbal cues for initiating opportunistic interactions. Although considerable research has aimed to enhance opportunistic interactions in video conferencing, this remains a challenging task.
2.3. Opportunistic Interactions in Social VR Events
@@
While VR has long been a key area of research in HCI (Burdea and Coiffet, 2003), its use in social contexts has only recently gained attention, driven by the rise of affordable commercial VR hardware, such as the Meta Quest [^2] Meta Quest: https://www.meta.com/quest/ and HTC Vive [^3] HTC Vive VR Headset: https://www.vive.com/us/. On social VR platforms, multiple users can interact with each other in 3D-rendered virtual spaces, accessible either via head-mounted displays (HMDs), or desktop setups using keyboard and mouse devices (McVeigh-Schultz et al., 2019, 2018).
Research has identified three major technical features unique to social interactions in VR. First, VR offers full-body avatars and real-time gesture capabilities within high-fidelity, immersive 3D environments (Freeman and Maloney, 2021). Second, it supports rich spatial and temporal experiences, fostering emotional connections similar to those in face-to-face interactions (Chang et al., 2024). Third, VR enables both verbal and embodied communication, allowing for a broad range of social activities (McVeigh-Schultz et al., 2019; Moustafa and Steed, 2018; Sra et al., 2018; Zamanifard and Freeman, 2019; Maloney et al., 2020b; Zhang et al., 2025). Consequently, unlike video conferencing, which is widely used in professional workplace settings, social VR platforms are naturally designed for more casual social gatherings (Maloney and Freeman, 2020), making them a promising venue for opportunistic interactions.
@@
For example, in a recent study, Freeman et al. interviewed 30 social VR users and found that their primary goal was to meet new people rather than maintain existing relationships. Participants built friendships through activities like dancing, sleeping, studying, homework help, and gaming, often connecting on first contact (Freeman and Acena, 2021). Further research by Maloney et al. observed how informal encounters in VR led to the initiation of opportunistic interactions. They found that a wide range of non-verbal signals were used to initiate connections, from simple nods and waves by avatars to more dynamic and sometimes “wild” gestures, such as dancing, kissing, pushing, bumping, and even blocking others’ view to capture attention (Maloney et al., 2020b). Many of these strategies would be uncommon or even considered inappropriate in face-to-face interactions between people meeting for the first time. Nevertheless, these “unconventional” behaviors are far more tolerated in the social VR context.
@@
More recently, an increasing number of professional events have been hosted in social VR, influenced by the COVID-19 pandemic (e.g.,  (Ahn et al., 2021; Moreira et al., 2022; Mulders and Zender, 2021)). For example, both the IEEE VR 2021 (Moreira et al., 2022) and VRARBB@SocialVR (Mulders and Zender, 2021) were academic conferences hosted on social VR platforms, Virbela and AltspaceVR, respectively. At IEEE VR 2021, attendees could choose between traditional tools such as YouTube and Twitch for talk broadcasting and text-based interaction, or immerse themselves in the Virbela space to watch talks and interact with other VR attendees. In VRARBB, participants engaged in topic-oriented discussions in AltspaceVR, spread across six weeks of conference time. While participants in both events reported high levels of engagement and found VR to be a suitable medium for academic gatherings, both studies provided limited detail on how social or spontaneous interactions actually unfolded in these environments. The above bulk of literature prompted our research question (RQ):
@@ 3.2. Event Contextualization
All 23 professional events in this study were public, stand-alone gatherings. Unlike academic conferences such as IEEE VR (Ahn et al., 2021; Moreira et al., 2022), which host a series of VR talks, panels, and discussions over multiple days with the same group of participants, the events we attended were one-off gatherings, primarily aimed at fostering networking opportunities among professionals within a time frame of 60 to 120 minutes. Attendees at these events usually shared a common interest, but they often did not know each other beforehand.

### v2

@@ Abstract.
social VR, opportunistic interactions, professional events, non-verbal cues
copyright: ccjournal: PACMHCIjournalyear: 2025journalvolume: 9journalnumber: 7article: CSCW519publicationmonth: 11doi: 10.1145/3757700ccs: Human-centered computing Collaborative and social computingccs: Human-centered computing Collaborative and social computing theory, concepts and paradigms
@@ 1. Introduction
The current research connects social VR with opportunistic interactions in professional settings—the unstructured yet essential exchanges that enable social bonding and knowledge transfer among members of the same workspace or community. Previous work in CSCW and communication research has delved deeply into each of these areas. Behavioral studies on how people coordinate spontaneity in workplace communication can date back to early 1990s, followed by two decades of experimental work aiming to reproduce this coordination in remote settings (e.g.,  (Tang and Rua, 1994; Fish et al., 1990; Isaacs and Tang, 1994)). Under a separate endeavor, recent scholarship has explored the emerging use of social VR for social hangouts and networking in informal meetups (e.g.,  (McVeigh-Schultz et al., 2019; Moustafa and Steed, 2018; Maloney and Freeman, 2020)). Little research, however, exists at the intersection of these two, although social VR arguably holds greater promise than most other mediums for supporting high-stakes interactions.
We offer an in-depth qualitative account to address this oversight. Specifically, we investigate factors that contribute to the successes and failures of opportunistic interactions in VR-mediated professional events. All participants involved in this investigation were authentic users of social VR. With their consent, we shadowed each person during one or more public professional events they have signed up for, closely observing the way they navigated each opportunistic interaction in its social context. We also conducted post-event interviews with them to gather their retrospections, enabling triangulation. The data obtained from these steps allow us to develop a nuanced understanding of how interaction unfolds in a context jointly characterized by spontaneity, professionalism, and virtuality.
Our findings provide rich empirical insights that solidify the promise and current roadblocks toward effectively initiating opportunistic interaction in social VR. At the center of these findings lies the notion of verisimilitude, or the participant’s perceived extent to which a social VR environment has preserved the real-world meanings of various non-verbal cues exchanged there. Our participants’ behaviors and interview responses revealed three distinct steps leading to the eventual occurrence of opportunistic interaction in professional settings: availability recognition, attention capture, and ice-breaking. At each step, an individual’s assessment of verisimilitude could be disrupted by a varied set of issues. Participants demonstrated an escalating need for high verisimilitude as they advanced through the three steps. They showed an increasing reliance on the host’s assistance when this need was not met. These insights provide actionable guidance for the design of future social VR platforms, as well as for people who navigate professional interactions via those platforms.
@@ 2. Related Work
This section summarizes prior studies and findings related to our current research. We begin by reviewing literature on opportunistic interactions in traditional face-to-face workplaces, emphasizing its crucial role in professional settings (Section 2.1). Next, we compare the similarities and differences between face-to-face and video-mediated opportunistic interactions for professional purposes (Section 2.2). We then synthesize existing insights on social VR as an emerging medium for opportunistic interactions, primarily in casual or recreational contexts. Together, these three strands of literature set the stage for our investigation of opportunistic interactions in professional social VR settings (Section 2.3).
2.1. Opportunistic Interactions in Professional Face-to-Face Settings
@@
Opportunistic interactions play a crucial role in professional face-to-face settings (Kraut et al., 1990; Fay, 2011). Studies show that in office environments, opportunistic interactions can constitute up to approximately 92% of all conversations (Whittaker et al., 1994), fostering strong social bonds among colleagues (Holmes and Marra, 2004), enabling spontaneous problem-solving without the formal pressures of meetings (Conrad and Poole, 1997), and helping to build meaningful interpersonal relationships (Duck, 1993). Beyond the workplace, opportunistic interactions is equally essential in professional contexts like conferences and workshops, where it facilitates networking, secures business clients and contracts, and promotes academic collaborations (Oester et al., 2017; McCarthy et al., 2004).
@@
Despite the prevalence of opportunistic interactions in professional settings, initiating these interactions can feel challenging due to the complex and dynamic behaviors involved (Pillet-Shore, 2018). Research in communication and psychology has therefore examined how face-to-face conversations are initiated, identifying two main features: audible and non-verbal cues (Kendon and Ferber, 1973; Pillet-Shore, 2012). For example, Pillet-Shore analyzed 337 video-recorded openings of face-to-face interactions, finding that people often use greeting utterances like hello or good morning to acknowledge another’s presence as a conversation starter. While these greetings may appear similar, they carry subtle nuances along a prosodic continuum, with “large” greetings expressing positivity and “small” ones conveying a neutral stance (Pillet-Shore, 2012).
@@
Non-verbal signals are also essential and often function in combination with audible cues to initiate opportunistic interactions in face-to-face settings. Previous research observed that visible smiling frequently accompanies greeting utterances, facilitating conversation openings (Kendon and Ferber, 1973). Additionally, individuals may adjust their body positions during moments of co-presence to ease the initiation process, allowing them to reset spatial and territorial orientations for a smoother approach (Goodwin, 1981; Mortensen and Hazel, 2014). In another study, Keevallik analyzed a full-day recording of 11 young people working together, finding that factors such as posture, gaze, and the bodily orientation of fellow workers contribute to attracting responses from others (Keevallik, 2018).
@@
In short, the above literature reveals that multiple elements, both verbal and non-verbal, contribute to the successful initiation of opportunistic interactions during face-to-face encounters. While multiple signals can jointly influence a person’s perceived availability and willingness to engage, the duration of this assessment is typically as brief as a few seconds (Pillet-Shore, 2018). The initiation of a face-to-face interaction is generally regarded as a single, cohesive, and brief process.
2.2. Video-Mediated Opportunistic Interactions in Professional Contexts
@@
Over recent decades, advancements in computer-mediated communication (CMC) tools have profoundly reshaped communication in professional contexts, transitioning from solely co-located, face-to-face interactions to more flexible, frequently remote environments (Fish et al., 1992; Zhao and Rosson, 2009; Egido, 1988). The shift was accelerated by the COVID-19 pandemic, which prompted many companies to adopt fully remote work models through team-based video conferencing (Fana et al., 2020; Bennett et al., 2021; Hacker et al., 2020). Despite this flexibility, remote settings greatly limit opportunistic interactions that naturally occur in physical workspaces (Fay, 2011; Bleakley et al., 2022). Previous literature attributes this reduction in CMC-based opportunistic interactions to two main factors: the lack of spatial relationships in video conferencing—similar to hallways in physical offices that facilitate casual encounters (Rudnicka et al., 2022; Isaacs and Tang, 1994)—and the absence of rich cues prevalent in face-to-face interactions (Grayson and Monk, 2003; Fish et al., 1992; Doherty-Sneddon et al., 1997).
@@
CSCW researchers have long investigated ways to foster a sense of co-location in distributed video conferencing (e.g.,  (Hauber et al., 2006; Raskar et al., 1998; Gibbs et al., 1999; Vertegaal, 1999; Benford et al., 2001; Elrod et al., 1992)). For example, early work by Fish et al. attempted to simulate physical proximity by placing human-scale screens in workspaces on different floors, encouraging workers in separate locations to engage in spontaneous conversations (Fish et al., 1990). Later, CRUISER (Fish et al., 1992; Root, 1988) and Montage (Tang et al., 1994; Tang and Rua, 1994) extended this concept to desktop-based video software. CRUISER was inspired by the concept of roaming a hallway and meeting people without prior planning. It facilitated both intentional and system-generated spontaneous video connections between users. Montage, on the other hand, introduced a gradual appearance and fading effect for the other person’s image, simulating movement closer to or farther away to mimic interactions in physical spaces.
While these early explorations were not directly integrated into later video conferencing systems, the concept of rendering proximity has become widely adopted. More recent commercial platforms such as Gather.town  [^1] Gather.town: https://www.gather.town/ use a shared 2D spatial layout where distributed users can see and approach each other. When users come into close proximity, a traditional video chat interface appears, allowing them to initiate spontaneous interactions. A recent study by Sanchez et al. found that such features indeed enhanced users’ awareness of each other, making it easier to determine when to interact, interrupt, or initiate a conversation. However, Gather.town users still reported insufficient information on whether others were actually interruptible for a conversation, essentially due to the lack of non-verbal cues within the 2D spatial layout interfaces (Palos-Sanchez et al., 2023).
@@
It should be noted that non-verbal cues, such as facial expressions, gaze, and gestures, are widely utilized during video conferencing (Nguyen and Canny, 2009). Studies have shown that gestures and facial expressions can convey subtle emotions (Isaacs and Tang, 1994), whereas gaze can foster trust when well-perceived through video streaming (Nguyen and Canny, 2007). However, unlike in face-to-face settings, these non-verbal exchanges in video-mediated interactions generally come into play only after a conversation has begun. The reliance on video cameras and 2D screens also limits the effectiveness of non-verbal cues for initiating opportunistic interactions (Sears et al., 2013). Thus, although considerable research has sought to enhance opportunistic interactions in video conferencing, it remains a challenging task.
2.3. Opportunistic Interactions in Social VR
@@
While VR has long been a key area of research in HCI (Burdea and Coiffet, 2024), its use in social contexts has only recently gained attention, driven by the rise of affordable commercial VR hardware, such as the Meta Quest [^2] Meta Quest: https://www.meta.com/quest/ and HTC Vive [^3] HTC Vive VR Headset: https://www.vive.com/us/. On social VR platforms, multiple users can interact with each other in 3D-rendered virtual spaces, accessible either via head-mounted displays (HMDs), or desktop setups using keyboard and mouse devices (McVeigh-Schultz et al., 2019, 2018).
@@
Research has identified three major technical features unique to social interactions in VR. First, VR offers full-body avatars and real-time gesture capabilities within high-fidelity, immersive 3D environments (Freeman and Maloney, 2020). Second, it supports rich spatial and temporal experiences, fostering emotional connections similar to those in face-to-face interactions (Chang et al., 2024). Third, VR enables both verbal and embodied communication, allowing for a broad range of social activities (McVeigh-Schultz et al., 2019; Moustafa and Steed, 2018; Sra et al., 2018; Zamanifard and Freeman, 2019; Maloney et al., 2020b; Zhang et al., 2025). Consequently, unlike video conferencing, which is widely used in professional workplace settings, VR has been naturally adopted for more casual social gatherings (Maloney and Freeman, 2020). In a recent study, Freeman et al. interviewed 30 social VR users and found that their primary goal was to meet new people rather than maintain existing relationships (Freeman and Acena, 2021). Further research by Maloney et al. observed how informal encounters in VR led to the initiation of opportunistic interactions. They found that a wide range of non-verbal signals were used to initiate connections, from simple nods and waves by avatars to more dynamic and sometimes “wild” gestures, such as dancing, kissing, pushing, bumping, and even blocking others’ view to capture attention (Maloney et al., 2020b). Many of these strategies would be uncommon or even considered inappropriate in face-to-face interactions between people meeting for the first time. Nevertheless, these “unconventional” behaviors are far more tolerated in the social VR context.
@@
More recently, a growing number of professional events have been hosted in social VR, triggered by the COVID-19 pandemic and continuing beyond it (e.g.,  (Ahn et al., 2021; Moreira et al., 2023; Mulders and Zender, 2021)). For example, IEEE VR 2021 allowed attendees to immerse themselves in the Virbela space to watch talks and interact with each other  (Moreira et al., 2023). Another academic event, VRARBB@SocialVR, leveraged AltspaceVR for topic-oriented discussions over six weeks of conference time  (Mulders and Zender, 2021). While participants in both events reported VR to be a suitable medium for professional gatherings, limited information is available on how opportunistic interactions actually unfolded in these environments. This body of literature prompted us to ask the following research question (RQ):
@@ 3.2. Event Contextualization
All 23 professional events in this study were public, stand-alone gatherings. Unlike academic conferences such as IEEE VR (Ahn et al., 2021; Moreira et al., 2023), which host a series of VR talks, panels, and discussions over multiple days with the same group of participants, the events we attended were one-off gatherings, primarily aimed at fostering networking opportunities among professionals within a time frame of 60 to 120 minutes. Attendees at these events usually shared a common interest, but they often did not know each other beforehand.